(d2y/dx2) - y = 4/(1+e^​x)


1
Expert's answer
2022-03-31T13:52:09-0400

At first, we solve the equation: "y''-y=0". To solve the equation, we substitute solution of the form:"g=e^{\lambda x}" and receive the equation for "\lambda:" "\lambda^2-1=0". Solutions are: "\lambda=\pm1". Thus, the solution of the homogenous equation "y''-y=0" is: "g=c_1e^{x}+c_2e^{-x}". The solution of the initial equation can be presented as: "y=g+f", where "f" is a solution of the initial equation. Substitute "f=ce^{x}" with function "c(x)" into the initial equation:

1). "c''e^{x}+2c'e^x+ce^x-ce^x=\frac{4}{1+e^x}". Do the change of variables: "f_1=c'" in the equation: "f_1'e^{x}+2f_1e^x=\frac{4}{1+e^x}" .Solve another homogeneous equation: "f_1'=-2f_1". The solution is: "f_1=b_1e^{-2x}". Suppose that "b_1" is a function. Substitute the latter expression into equation "f_1'e^{x}+2f_1e^x=\frac{4}{1+e^x}" and receive: "b_1'=\frac{4e^{x}}{1+e^x}" Compute the integral to find "f_1": "b_1=4\int\frac{d(e^x)}{1+e^x}=4(ln(1+e^x)+C_1)". "f_1=4e^{-2x}ln(1+e^x)+e^{-2x}C_1". "c=\int fdx=\int(4e^{-2x}ln(1+e^x)+4e^{-2x}C_1)dx=I_1+I_2", where "I_1=4\int e^{-2x}ln(1+e^x)dx", "I_2=4\int e^{-2x}C_1dx". Compute the integral "I_1" :

"I_1=4\int e^{-2x}ln(1+e^x)dx". Make the change of variables: "z=e^{x}". The integral becomes: "I_1=4\int z^{-3}ln(1+z)dz=-2{z^{-2}}ln(1+z)+2\int \frac{z^{-2}}{1+z}dz=-2{e^{-2x}}ln(1+e^x)+2(-e^{-x}+ln(1+e^x)-x)+a_1"

"I_2=-2C_1e^{-2x}+a_2"

Thus, "c=-2{e^{-2x}}ln(1+e^x)+2(-e^{-x}+ln(1+e^x)-x)+a_1-2C_1e^{-2x}+a_2" and, respectively:

"f=-2+2(e^{x}-e^{-x})ln(1+e^{x})-2xe^x-2C_1e^{-x}+Ke^x". Last two terms can be absorbed by "g" and "y=c_1e^{x}+c_2e^{-x}-2+2(e^{x}-e^{-x})ln(1+e^{x})-2xe^x" This is the solution of the problem.